OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5x + 14 = 26

OpenStudy (anonymous):

subtract 14 on both sides then divide 5 on both sides for your answer

OpenStudy (anonymous):

sorry. 5^x

OpenStudy (anonymous):

subtract 14 both sides. use natural log on both sides. bring down x. divide ln5 on both sides. you get x

OpenStudy (anonymous):

You got it? you could also do.. when you subtract 14 both sides.. use logx a = b which is also x^b = a

OpenStudy (anonymous):

no im kind of confused

OpenStudy (anonymous):

...hm... its 5^x + 14 = 26, right? show me what you got so far from what is mentioned above.

OpenStudy (anonymous):

log5^x=log12 right?

OpenStudy (anonymous):

I got ln5^x=2.48

OpenStudy (anonymous):

yes. that is correct. but I see you are not in the natural log yet. 5^x + 14 = 26 5^x = 12 log(base 5) 12 = x if you do ln... ln5^x = ln12 xln5 = ln 12 x = ln12/ln5 x = log(base 5) 12

OpenStudy (anonymous):

ln5^x = 2.48 is right but that is not your purpose. what you want is x. and so, that is what i mentioned.

OpenStudy (anonymous):

okay. well after i enter it into my calculator i still have the wrong answer.

OpenStudy (anonymous):

what did you enter in your calculator?

OpenStudy (anonymous):

log(5)*12

OpenStudy (anonymous):

log doesn't work that way. try: log(12)/log(5) <---- 5 would be considered the base. then try: 5^(log(12)/log(5)) + 14

OpenStudy (anonymous):

because when you type log(n) in the calculator, its in base 10.

OpenStudy (anonymous):

i thought that's why i put log(5) doesn't that make 5 the base?...

OpenStudy (anonymous):

no... when you type in: log(5) <-- the calculator would treat this as log(5)/log(10) <---the base is 10 and so when you type log(12)/log(5): the calculator would see this as: ( log(12) / log(10) ) / ( log(5) / log(10) ) the log(10)'s would cancel.

OpenStudy (anonymous):

okay i get it now

OpenStudy (anonymous):

Ok thats good.

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

:) no probs. you did understand how we got the x, right? not just the calculator part.

OpenStudy (anonymous):

yeah i got it. i just did another one :)

OpenStudy (anonymous):

Ok cool.